The solubility of CO2 in water at 25°C and 1 atm is 0.034 mol/L. What is its solubility at a partial pressure in air of 0.00049 atm? Assume that CO2 obeys Henry's law.
Answer #1

Henry's law :

P = KHC

KH = P / C

KH = 1 / 0.034

= 29.4 Atm Litre/mol

Now the partial pressure is 0.00049 atm , so P = 0.00049 atm

C = P / KH

C = 0.00049 / 29.4

C = 1.67 x 10-5 M